Answer the questions below.

(a) How many different committees of size

3

can be formed from

15

people?
(b) There are

17

appetizers available at a restaurant. From these, John is to choose

13

for his party. How many groups of

13

appetizers are possible?
Answer #1

1. The required number of committees can be formed in 15C3 ways , i.e \frac{15!}{3!*12!}

= 13*14*15/6 = 455

2. This can be done in 17C13 = \frac{17!}{13!*4!}

= 2380 ways